OpenStudy (anonymous):

Based on the quadratic formula, explain the solution to a quadratic equation if the value of b² is smaller than the value of 4ac.

OpenStudy (goldphenoix):

Well, the quadratic formula is  Let say a quadratic equation is 3x^2+10x-25=0. b = 10, a = 3, c = -25. b^2 = 10^2=100 4ac=4(3)(-25)=-300 x = 5/3, -5 As shown, b^2 is larger than 4ac. If b^2 is smaller than 4ac, then the value will be irrational, or an imaginary number. Why? Because if a smaller is subtracted by a larger number in a radical, then the value will be irrational. For instance, b= 5, 4ac = 10 \[\sqrt{5-10} = \sqrt{-5}\]